OpenStudy (mven):

A square pyramid has sides of 8 m. Each face is an equilateral triangle. What is the lateral area of the pyramid?

OpenStudy (anonymous):

The area of an equilateral triangle is s^2√(3)/4, where s is a side length. The lateral area of this square pyramid is made up of a 4 equilateral triangles, so the lateral area is 4s^2√(3)/4 or just s^2√3. LA = s^2√3 = 8^2√3 = 64√3 = 110.85 m^2 ANSWER: about 111 m^2

OpenStudy (mven):

Here is what i get

OpenStudy (anonymous):

Hmm hold on..

OpenStudy (mven):

haha thanks

OpenStudy (mven):

@Kyliejones any idea??